Suppose the current exchange rate for the Polish zloty is Z 2.18. The expected exchange rate in three years is Z 2.55. What is the difference in the annual inflation rates for the United States and Poland over this period? Assume that the anticipated rate is constant for both countries.
Answer #1

Let inflation rate in US be us and in Poland be pol

(1+us)^3=2.18*(1+pol)^3/2.55

=>(1+us)/(1+pol)=(2.18/2.55)^(1/3)

=>(1+us)/(1+pol)=0.9491

=>us-pol=-5.09%